# Structural Alignment Improves Graph Test-Time Adaptation

### v1

@@ 6 Discussion and Limitations
In this work, we theoretically analyze the impact of various graph shifts on the generalization gap in GTTA and highlight their empirical effects on test-time performance degradation. Based on these insights, we propose the TSA algorithm, which addresses neighborhood shift through extended and improved neighborhood alignment strategies tailored to the unique challenges of GTTA, and handles label shift via classifier adjustments.
Despite its effectiveness, we observe that TSA ’s benefits can be marginal when feature shift dominates, as evidenced in the last two columns of Pileup, Arxiv, and DBLP/ACM. This is primarily because the neighborhood alignment strategy is not robust in such scenarios. Moreover, our current estimation of $\bm{\gamma}$ heavily depends on the accuracy of refined soft pseudo-labels. While confusion-matrix-based methods (Lipton et al., 2018; Azizzadenesheli et al., 2019; Alexandari et al., 2020) can provide more robust estimations, they often require additional adversarial training, which is not compatible with GTTA constraints. We encourage future work to explore more robust approaches for estimating $\bm{\gamma}$ to further enhance adaptation under severe feature shifts.
7 Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Omitted Proofs
The function of mean pooling naturally marginalizes the effect of degree magnitude, meaning that only the expectation of the neighboring nodes’ label distribution and the label distribution of the central node itself influence the prediction. We use Lemma A.1 to bound the terms above. Since $\forall j\in\mathcal{Y},\gamma^{\mathcal{S}}_{u_{j}}\text{ and }\gamma^{\mathcal{T}}_{u_{j}}\in[0,1]$ and we have $\alpha_{j}+\beta_{j}=1$, we obtain:
@@
(a) comes from Lemma A.1. (b) comes from setting $\alpha_{j}=1$ and $\beta_{j}=0,\forall j\in\mathcal{Y}$ (c) The inequality comes from Holder’s inequality, where we take the maximum-norm for $k$ and take the $L_{1}$ norm to marginalize $Y_{v}$: $\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)=\sum_{k\in\mathcal{Y}}\mathbb{P}^{\mathcal{U}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ (d) The last term in the previous line corresponds to the conditional error gap in (Tachet des Combes et al., 2020) but in a GNN mean pooling fashion.
The term $\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ can be interpreted as the probability of error prediction using a mean pooling GNN encoder. By definition, $Y_{v}$ represents a random variable of the neighborhood label ratio and $Y_{u}$ denotes a random variable of self class label, their realization corresponds to the mean pooled and center value. As a reult, $\Delta_{CE}$ only exists if and only if feature shift exists. Putting everything together, we have
@@
$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2TV(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Neighborhood shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature shift}}\vspace{-2mm}$
Appendix B Additional Results
B.1 Results of GPRGNN
Table 5 presents the results of the GPRGNN backbone on the MAG dataset. Compared to GraphSAGE, we observe that GPRGNN is less robust under structure shifts due to its linear combination of k-hop representations. We also notice a similar tendency as with GraphSAGE, where T3A outperforms all other baselines. TSA slightly improves upon T3A by refining the decision boundary. However, we frequently observe a performance decrease when combining AdaRC with T3A. This drop is attributed to the fact that their proposed PIC loss relies on minimizing intra-variance, which depends on the quality of node hidden representations on the target domain. Under imbalanced training and label shift, node hidden representations can easily degrade, causing AdaRC performance to become unstable.
@@ B.1 Results of GPRGNN
Table 5: MAG GPRGNN results (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
@@
AdaRC-TENT 13.90$\pm$0.71 25.87$\pm$0.47 29.30$\pm$0.82 18.14$\pm$0.19 21.03$\pm$0.26 16.38$\pm$0.80 7.62$\pm$0.73 13.84$\pm$0.74 7.85$\pm$0.81 5.97$\pm$0.36
AdaRC-LAME 8.93$\pm$4.76 8.18$\pm$0.78 15.46$\pm$1.34 6.25$\pm$1.43 6.69$\pm$0.81 1.35$\pm$2.69 17.60$\pm$8.89 33.42$\pm$2.41 18.81$\pm$1.42 4.23$\pm$8.02
AdaRC-T3A 22.34$\pm$18.21 32.18$\pm$1.70 42.85$\pm$1.15 34.77$\pm$2.62 25.75$\pm$0.62 45.27$\pm$2.57 33.01$\pm$1.17 42.08$\pm$0.82 38.57$\pm$1.50 29.30$\pm$0.47
TSA-TENT 14.25$\pm$0.68 25.76$\pm$0.51 29.30$\pm$0.79 18.30$\pm$0.26 21.20$\pm$0.20 16.57$\pm$0.78 7.69$\pm$0.72 14.03$\pm$0.70 8.07$\pm$0.79 6.03$\pm$0.39
TSA-LAME 26.33$\pm$2.25 22.76$\pm$1.44 35.57$\pm$1.43 18.15$\pm$2.08 16.94$\pm$1.06 29.70$\pm$0.45 11.04$\pm$0.61 17.90$\pm$0.89 12.46$\pm$0.43 9.32$\pm$0.51
TSA-T3A 40.81$\pm$0.23 39.13$\pm$0.77 49.16$\pm$0.85 47.05$\pm$1.74 33.27$\pm$1.21 41.06$\pm$3.34 38.34$\pm$1.25 46.38$\pm$1.83 43.64$\pm$0.85 28.71$\pm$0.57
B.2 Results of Ablation Study
Table 6: Ablation Study on TSA.

### v2

@@ 6 Conclusion
In this work, we theoretically analyze the impact of various graph shifts on the generalization gap in GTTA and highlight their empirical effects on test-time performance degradation. Based on these insights, we propose TSA, a method that addresses structure shift through neighborhood alignment, SNR-inspired adjustment, and boundary refinement. Extensive experiments validate the effectiveness and efficiency of TSA.
@@ Appendix A Omitted Proofs
Most GNNs marginalize the effect of degree magnitude, meaning that only the expectation of the neighboring nodes’ label distribution and the label distribution of the central node itself influence the prediction. We use Lemma A.1 to bound the terms above. Since $\forall j\in\mathcal{Y},\gamma^{\mathcal{S}}_{u_{j}}\text{ and }\gamma^{\mathcal{T}}_{u_{j}}\in[0,1]$ and we have $\alpha_{j}+\beta_{j}=1$, we obtain:
@@
(a) comes from Lemma A.1. (b) comes from setting $\alpha_{j}=1$ and $\beta_{j}=0,\forall j\in\mathcal{Y}$ (c) The inequality comes from Holder’s inequality, where we take the maximum-norm for $k$ and take the $L_{1}$ norm to marginalize $Y_{v}$: $\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)=\sum_{k\in\mathcal{Y}}\mathbb{P}^{\mathcal{U}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ (d) The last term in the previous line corresponds to the conditional error gap in [58] but in a standard GNN’s message passing fashion.
The term $\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ can be interpreted as the probability of error prediction using a GNN encoder. By definition, $Y_{v}$ represents a random variable of the neighborhood label ratio and $Y_{u}$ denotes a random variable of self class label, their realization corresponds to the expectation of the neighboring nodes’ label distribution and the label distribution of the central node itself. As a result, $\Delta_{CE}$ only exists if and only if feature shift exists. Putting everything together, we have
@@
$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2TV(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Conditional Structure Shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature Shift}}\vspace{-2mm}$
Appendix B Limitations
While TSA can address feature shift to some extent, we observe that TSA ’s benefits can be marginal when feature shift dominates significantly over structure shift, as evidenced in the last two columns of Pileup. This is primarily because the neighborhood alignment strategy is not robust in such scenarios. Moreover, our current estimation of $\bm{\gamma}$ heavily depends on the accuracy of refined soft pseudo-labels. While confusion-matrix-based methods [72, 73, 74] can provide more robust estimations, they often require additional adversarial training, which is not compatible with GTTA constraints. We encourage future work to explore more robust approaches to estimating $\bm{\gamma}$ to further enhance adaptation under severe feature shifts.
Appendix C Broader Impact
Our work studies structure shift and proposes TSA for graph test-time adaptation (GTTA). The goal of GTTA is to enhance model performance under distribution shifts with computation and privacy constraints. We do not foresee any direct negative societal impacts.
Appendix D Additional Results
D.1 TSA Ablation Studies
In the ablation studies, we examine the effects of neighborhood alignment and SNR-based representation adjustments. We select synthetic datasets with structure shift, as well as real-world datasets that exhibit the most conditional structure shift (CSS), according to Table LABEL:table:magstats. A detailed discussion of the ablation results is provided in Section 5.2.
Table 5: TSA Ablation studies on CSBM and MAG.
[…]
TSA-TENT Base 88.68$\pm$4.99 66.25$\pm$7.75 32.53$\pm$0.91 28.62$\pm$0.63
w/o Nbr. Align 78.45$\pm$5.51 61.12$\pm$4.86 32.26$\pm$0.95 28.41$\pm$0.59
w/o SNR 86.93$\pm$5.05 63.82$\pm$8.06 32.52$\pm$0.90 28.57$\pm$0.59
TSA-LAME Base 65.09$\pm$2.34 52.89$\pm$6.11 32.86$\pm$2.23 27.22$\pm$1.48
w/o Nbr. Align 48.45$\pm$5.30 37.31$\pm$3.95 32.70$\pm$2.29 26.49$\pm$1.76
w/o SNR 63.65$\pm$2.65 51.18$\pm$5.84 29.89$\pm$1.94 24.41$\pm$1.43
TSA-T3A Base 65.59$\pm$2.57 52.34$\pm$7.19 46.61$\pm$0.88 43.45$\pm$0.81
w/o Nbr. Align 45.69$\pm$5.44 36.75$\pm$3.09 46.44$\pm$0.80 42.09$\pm$1.26
w/o SNR 64.17$\pm$2.11 51.04$\pm$6.79 46.51$\pm$0.94 41.35$\pm$1.09
D.2 Visualization of Distribution Shift
Figure 5 (a)-(e) investigates the impact of different types of distribution shifts on GNNs. We use a one-layer GraphSAGE followed by an MLP classifier to visualize node representations on synthetic datasets generated by the CSBM model.
  • Feature Shift: Figure 5 (b) shows that the representations deviate from the source domain, but the clusters remain separable.
  • Conditional Structure Shift (CSS): Figure 5 (c) shows that the embeddings largely overlap, making the representation space less discriminative.
  • Label Shift: Figure 5 (d) illustrates a biased decision boundary, which leads to misclassification of the minority class.
  • SNR Shift: Figure 5 (e) shows that the node representations remain roughly aligned with the source, but are more dispersed, as the SNR shift introduces noise into the representations.
(a) Source (b) Feature Shift (c) CSS (d) Label Shift (e) SNR Shift
Figure 5: The t-SNE visualization of a one-layer GraphSAGE under different distribution shifts: (top) the output representations given by the GNN encoder and (bottom) the node representations with the classifier decision boundaries. (a) indicates the source domain where the model is pretrained. (b), (c), and (d) stand for feature shift, conditional structure shift (CSS), and label shift. (e) represent the impact of SNR shift induced by degree changes. The color of the nodes represents the ground-truth labels. The legends in the bottom plots show the accuracy and recall scores for each class. The red contours in the bottom (d) highlight the decision boundary of the minority class.
Table 6: Computation time on MAG (CN $\rightarrow$ US). TTA methods are recorded in adaptation time per epoch. The Wall-Clock Time (seconds) are measured on a single NVIDIA RTX A6000 GPU.
Method Stage Computation Time (sec) Additional Time (%) Inference 0.518$\pm$0.038 TSA-T3A Adaptation 1.138$\pm$0.072 219.69% - Neighborhood Alignment 0.037$\pm$0.011 7.14% - SNR Adjustment 0.729$\pm$0.023 140.73% - Boundary Refinement (T3A) 0.368$\pm$0.046 71.04% GTrans Adaptation 4.309$\pm$0.171 831.85% HomoTTT Adaptation 2.014$\pm$0.815 388.80% SOGA Adaptation 9.264$\pm$0.218 1788.42%
D.3 Computational Efficiency
We quantify the computation time of TSA when adapting GraphSAGE on the MAG US dataset and compare it with other GTTA methods:
  • Inference time refers to the time required for the model to generate predictions on the target graph ${\mathcal{G}}^{{\mathcal{T}}}$. This is also equivalent to the computation time for ERM.
  • Adaptation time refers to the time required per epoch after inference. It reflects the additional overhead introduced by the adaptation methods.
For a fair comparison, we report the computation time per epoch of adaptation. Note that GTrans alternates between optimizing adjacency structure and node features; in this case, we report the combined runtime of one graph feature update and one edge update as a single adaptation epoch. Table 6 shows that TSA is more efficient than other GTTA baselines. A detailed discussion of computational efficiency is provided in Section 5.2.
D.4 Hyperparameter sensitivity
Hyperparameter analysis on $\rho_{1}$. TSA employs neighborhood alignment based on the uncertainty of target pseudo labels. To evaluate the sensitivity of the hyperparameter $\rho_{1}$ (as defined in Section 4.1 by the entropy threshold $-\sum_{i\in{\mathcal{Y}}}[\hat{y}]_{i}\ln([\hat{y}]_{i})\leq\rho_{1}\cdot\ln(|{\mathcal{Y}}|)$), we vary $\rho_{1}$ within the range of $[-0.1,0.1]$ centered around the value selected on the validation set. We perform evaluations on both synthetic and real-world datasets, using the best-performing variants: TSA-TENT, TSA-T3A and TSA-LAME for each case. TSA-TENT and TSA-T3A are evaluated with the GraphSAGE backbone, while TSA-LAME is evaluated with the GPRGNN backbone.
Overall, Figure 6 shows that TSA is more robust to variations in $\rho_{1}$ on synthetic datasets, but more sensitive on real-world datasets due to the greater task complexity. Nevertheless, all tested values of $\rho_{1}$ outperform the corresponding non-graph TTA baselines.
Hyperparameter analysis on $\rho_{2}$. When optimizing $\alpha$ using the cross-entropy loss in Eq. 5, real-world datasets often produce unreliable hard pseudo-labels. To address this, we adopt the same strategy used for selecting $\rho_{1}$. Only hard pseudo-labels whose entropy satisfies $-\sum_{i\in{\mathcal{Y}}}[\hat{y}]_{i}\ln([\hat{y}]_{i})\leq\rho_{2}\cdot\ln(|{\mathcal{Y}}|)$ are used to optimize $\alpha$. We select $\rho_{2}$ from $\{0.1,1.0\}$ for real-world datasets and fix $\rho_{2}=1$ for synthetic datasets. To further analyze the effect of $\rho_{2}$, we vary it within the range $[-0.1,0.1]$ centered around the value selected on the validation set.
Figure 7 shows that  TSA is robust to variations in $\rho_{2}$, with all perturbations outperforming the corresponding non-graph TTA baselines.
(a) CSBM (Struct. Shift) (b) MAG (US $\rightarrow$ FR) (c) Pileup (PU10$\rightarrow$30)
Refer to caption Refer to caption Refer to caption
Figure 6: Hyperparameter sensitivity analysis of $\rho_{1}$.
(a) MAG (US $\rightarrow$ FR) (b) Pileup (PU10$\rightarrow$30)
Refer to caption Refer to caption
Figure 7: Hyperparameter sensitivity analysis of $\rho_{2}$ on real-world datasets.
@@ D.4 Hyperparameter sensitivity
Table 7: GPRGNN results on MAG (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
@@
Matcha-TENT 13.90$\pm$0.71 25.87$\pm$0.47 29.30$\pm$0.82 18.14$\pm$0.19 21.03$\pm$0.26 16.38$\pm$0.80 7.62$\pm$0.73 13.84$\pm$0.74 7.85$\pm$0.81 5.97$\pm$0.36
Matcha-LAME 8.93$\pm$4.76 8.18$\pm$0.78 15.46$\pm$1.34 6.25$\pm$1.43 6.69$\pm$0.81 1.35$\pm$2.69 17.60$\pm$8.89 33.42$\pm$2.41 18.81$\pm$1.42 4.23$\pm$8.02
Matcha-T3A 22.34$\pm$18.21 32.18$\pm$1.70 42.85$\pm$1.15 34.77$\pm$2.62 25.75$\pm$0.62 45.27$\pm$2.57 33.01$\pm$1.17 42.08$\pm$0.82 38.57$\pm$1.50 29.30$\pm$0.47
TSA-TENT 14.51$\pm$0.66 25.86$\pm$0.54 29.47$\pm$0.73 18.76$\pm$0.39 21.48$\pm$0.23 16.96$\pm$0.74 7.89$\pm$0.72 14.32$\pm$0.70 8.35$\pm$0.81 6.16$\pm$0.38
TSA-LAME 26.58$\pm$2.09 22.75$\pm$0.85 35.95$\pm$0.98 19.30$\pm$1.99 17.49$\pm$0.79 32.30$\pm$0.67 14.86$\pm$1.12 21.88$\pm$0.95 15.02$\pm$0.86 11.13$\pm$0.51
TSA-T3A 40.77$\pm$0.15 39.03$\pm$0.86 49.38$\pm$0.89 47.81$\pm$1.81 32.43$\pm$1.24 40.94$\pm$2.82 38.12$\pm$1.42 46.26$\pm$1.81 43.70$\pm$1.33 29.07$\pm$0.46
Table 8: GPRGNN results on Pileup (f1-scores). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.